Consider the function P(x)=x(30-x). Is this function odd, even or neither? What do you expect if you graph this function?
Answer by stanbon(48568)

P(x)=x(30-x) = 30x-x^2
P(-x) = -x(30+x) = -30x-x^2
-P(-x) = 30x+x^2
=============
Since P(x) does not equal P(-x) P(x) is not even
Since P(x) does not equal -P(-x) P(x) is not odd
------------------
The graph will not be symmetric to the y axis or to the origin.
=================
Cheers,
Stan H.